# Compressing local vertex functions from the multipoint numerical renormalization group using quantics tensor cross interpolation

### v2

@@ II.2 Matsubara formalism
$K(\boldsymbol{\Omega}_{p})=\!\begin{cases}&\prod_{i=1}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}^{-1}\hfill\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \prod_{i=1}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}\neq 0,\\[8.5359pt] &\!-\frac{1}{2}\!\biggl{[}\beta+\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}^{-1}\biggr{]}\!\prod\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}^{-1}\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \exists j:\Omega_{\overline{1}\cdots\overline{j}}=0,\end{cases}$ (7)
@@
$G_{p}(\mathrm{i}\boldsymbol{\omega}_{p})=\sum_{{\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}}}S[\boldsymbol{\mathcal{O}}_{p}]({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}})\prod_{i=1}^{3}(\mathrm{i}{\omega}_{\bar{1}\cdots\bar{i}}-{\epsilon}_{\bar{1}\cdots\bar{i}})^{-1}{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}{.}}$ (8)
@@ II.3 Keldysh formalism
$K_{b}^{[\lambda]}(\boldsymbol{\omega}_{p},\boldsymbol{\epsilon}_{p})$ $=\prod_{j=1}^{\ell-1}\lim_{\gamma_{0}\rightarrow 0^{+}}\int_{\mathbb{R}}\mathrm{d}\omega_{\overline{1}\ldots\overline{j}}^{\prime}\,\frac{\delta_{b}(\omega_{\overline{1}\ldots\overline{j}}^{\prime},\epsilon_{\overline{1}\ldots\overline{j}})}{\omega_{\overline{1}\ldots\overline{j}}-\omega_{\overline{1}\ldots\overline{j}}^{\prime}+\mathrm{i}\gamma^{\lambda}_{0,j}}\,{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}{,}}$ , (10b)
@@
$\mathrm{i}\gamma^{\lambda}_{0,j}$ =$=\begin{cases}\mathrm{i}\gamma_{0}\cdot(\ell-j)&j\geq\lambda,\\ -\mathrm{i}\gamma_{0}\cdot j&j<\lambda.\\ \end{cases}$ (10c)
@@ II.4 Symmetric improved estimators: from correlators to the vertex
is a $2\!\times\!2$ matrix acting on the $i$th Keldysh index of $G_{\mathrm{con}}$. In the Matsubara formalism, $X$ is replaced by scalar unity. In practice, the main workload in computing $\Gamma_{\mathrm{core}}$ at a given frequency $\boldsymbol{\omega}$ is the evaluation of all $162^{4}=$ 16 connected correlators, each comprised of $4!=24$ partial correlators. The quantities $\mathcal{K}_{1}^{r}$ and $\mathcal{K}_{2^{(\prime)}}^{r}$ can be computed using an analogous approach presented in App. B.
@@ II.5 Quantics Tensor Cross Interpolation
Here, the $\alpha_{\ell}$ are virtual bond indices with $\ell$-dependent bond dimensions, $\alpha_{\ell}=1,\kern-0.74997pt.\kern-0.50003pt.\kern-0.50003pt.,\chi_{\ell}$. The maximum bond dimension, $\chi=\max{\chi_{\ell}}$, is called the rank of $F_{\boldsymbol{\sigma}}^{\mathrm{QTCI}}$. The maximum in Eq. (19) is estimated using all sampled entries of $F_{\boldsymbol{\sigma}}$. The TCI algorithm optimizes the tensors $[M_{\ell}^{\sigma_{\ell}}]_{\alpha_{\ell-1}\alpha_{\ell}}$ iteratively, progressively sampling $F_{\boldsymbol{\sigma}}$, until no $\boldsymbol{\sigma}$ is found where the error $\varepsilon_{\boldsymbol{\sigma}}$ exceeds a given tolerance $\tau$. During this process, the bond dimensions $\chi_{\ell}$ are increased dynamically to improve the accuracy of the tensor train representation. Finding a tensor train representation of $F^{\mathrm{QTCI}}_{\boldsymbol{\sigma}}$ with TCI has a computational cost of ($\order{R\chi^{3}}$ ).
@@ II.6 Implementation details
To obtain the core and full vertices in the Matsubara and Keldysh formalisms as QTTs, we apply QTCI to functions that evaluate $\Gamma_{\mathrm{core}}(\boldsymbol{\omega})$ and $\Gamma(\boldsymbol{\omega})$ on individual frequency points $\boldsymbol{\omega}$ to be specified on demand by the TCI algorithm. The frequencies $\boldsymbol{\omega}$ reside on an equidistant grid of $2^{3R}$ points. For Matsubara grids, the grid spacing is set by the temperature, and the extent of the grid can be increased exponentially by increasing $R$. For Keldysh vertices, which are functions of continuous frequencies, one may exponentially increase either the density of grid points, or the extent of the grid, or both, by increasing $R$. It is important not to precompute the vertices on a dense grid, as this precomputation step would incur costs scaling as ($\order{2^{3R}}$ ). By avoiding precomputation, $R$ can be increased to yield grid sizes and/or grid densities beyond those attainable by conventional means. The TCI algorithm samples a sparse set of ($\order{\chi^{2}R}$ ) points $\boldsymbol{\omega}$, which is much smaller than $2^{3R}$, the total number of grid points for the 4p vertex functions considered in this work. For this application, function evaluation during sampling is the dominant cost as opposed to the ($\order{\chi^{3}R}$ ) cost of computing prrLU factorizations (see [38, Sec. 3.3]). More specifically, the computational effort is dominated by the evaluation of partial 4p correlators. These enter $\Gamma_{\mathrm{core}}$ and $\Gamma$ via the full correlators appearing in Eq. (II.4). In this section, we discuss how our code evaluates partial correlators in an efficient way. Readers only interested in our results on the compressibility of Matsubara and Keldysh vertices can move on to Sec. III.
@@ II.6.1 Matsubara vertices
$G_{p}(\mathrm{i}\boldsymbol{\omega}_{p})=\!\sum_{{\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}}}k_{\omega_{\bar{1}}\epsilon_{\bar{1}}}k_{\omega_{\bar{1}\bar{2}}\epsilon_{\bar{1}\bar{2}}}k_{\omega_{\bar{1}\bar{2}\bar{3}}\epsilon_{\bar{1}\bar{2}\bar{3}}}S[\boldsymbol{\mathcal{O}}_{p}]({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}}).$ (20)
Since $({\omega}_{\bar{1}},{\omega}_{\bar{1}\bar{2}},{\omega}_{\bar{1}\bar{2}\bar{3}})$ and $({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}})$ live on finite frequency grids, the kernels $k_{{\omega}_{\bar{1}\cdots\bar{i}}{\epsilon}_{\bar{1}\cdots\bar{i}}}$ can be viewed as matrices. A typical grid size for the spectral function peaks $({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}})$ is $70\times 70\times 70$, while the largest grids for Matsubara frequencies $({\omega}_{\bar{1}},{\omega}_{\bar{1}\bar{2}},{\omega}_{\bar{1}\bar{2}\bar{3}})$ used in this work have $40962^{12}=$ 4096 points in each dimension. Thus $k_{{\omega}_{\bar{1}\cdots\bar{i}}{\epsilon}_{\bar{1}\cdots\bar{i}}}$ can be precomputed and stored for all relevant grid sizes.
@@
(i) Compressing the kernels is the more general of the two methods, in that it has a smaller memory footprint ($<1$ GB per full correlator $G(\mathrm{i}\boldsymbol{\omega})$ for $R=12,\tau=10^{-3}$). The idea is to compress the kernels $k_{\omega\epsilon}$ by exploiting their low-rank structure [39, 40]: We SVD-decompose each $k_{\omega\epsilon}$ and discard singular values below a given cutoff $S_{\mathrm{cut}}$, resulting in the approximation
@@
$\left|\sum_{\sum_{i}a_{i}>N}A_{a_{1}a_{2}a_{3}}\prod_{i=1}^{3}U_{\omega_{i}a_{i}}\right|$
$\qquad\leq\sqrt{\sum_{\sum_{i}a_{i}>N}|A_{a_{1}a_{2}a_{3}}|^{2}}\,\prod_{i=1}^{3}\max_{\omega_{i}}\sqrt{\sum_{a_{i}}\left|U_{\omega_{i}a_{i}}\right|^{2}}$
$\qquad=\sqrt{\sum_{\sum_{i}a_{i}>N}|A_{a_{1}a_{2}a_{3}}|^{2}}\;.$ (24)
@@
In this approach, we only have two kernel contractions in each evaluation $G_{p}$, but have to store the intemediates (26) for all partial correlators. Their size grows linearly in the grid frequency grid size, i.e., as $2^{R}$ where $R$ is the number of quantics bits. For example, if $R=12$ and the PSFs live on a $70\!\times\!70\!\times\!$ 70 × 70 logarithmic grid (which results from $2\times 6$ decades of energy bins with 8 points per decade and discarding zeros in the PSFs), each full correlator consumes $12.9$ GB of memory. On an $R=12$ grid at $\beta=2000,\,u=0.5$, the precomputation also yields a speedup of about a factor 60. But in contrast to the compression of kernels (Eq. (21)), this speedup is independent of the TCI tolerance $\tau$. Overall, method (ii) is recommended as long as its memory demands can be met, because it evaluates correlators in a numerically exact fashion.
@@ II.6.2 Keldyh vertices
We tested our code for evaluating vertex functions with the sIE scheme against the prior Matlab implementation used in Ref. 15. We found numerically exact agreement with a normalized discrepancy $<10^{-13}$ for the Matsubara quantities. In Keldysh the maximum discrepancy in the core vertex between our Julia code and the Matlab code of Ref. 15 is about $0.002\cdot||\Gamma_{\mathrm{core}}||_{\infty}$ (with the supremum norm $||\cdot||_{\infty}$). This discrepancy can be attributed to small differences in the broadening implementation, mainly the interpolation of the broadened kernel from a logarithmic to a linear grid (cf. App. A). This discrepancy is one order of magnitude smaller than the error introduced by the arbitrariness inherent in the choice of broadening parameters. As an additional test, our code was used to generate Keldysh vertex data to check the fulfillment of exact diagrammatic relations of mpNRG data [16].
@@
Figure 2: Singular values $S_{i}$ of regular Matsubara kernels (Eq. (7), solid line) and broadened, fully retarded Keldysh kernels (Eq. (28a), dashed line). We show kernels at inverse temperatures $\beta\in\{20,200,2000\}$ and interaction $u=0.5$. The frequency grids are bosonic with $2^{12}$ points, with the Keldysh grid ranging from $-0.65$ to $\omega_{\max}=0.65$.
@@ III.1 Single impurity Anderson model
Let us first consider the QTCI-compression of the Matsubara core and full vertices. An important input to the QTCI algorithm is the specified error tolerance $\tau$, see Eq. (19). When compressing vertices from mpNRG, the choice of tolerance should be based on the accuracy of the PSFs. Based on benchmark results of Refs. 11, 12, 15, we expect the mpNRG vertex to be reliable to roughly $10^{-3}\!\cdot\!||\Gamma_{\mathrm{core}}||_{\infty}$, where the error is partially systematic (as opposed to pure white noise). It should be emphasized that this is only an estimate and inherent errors in mpNRG (due to discretization of the noninteracting bath and discarding high-energy eigenstates during iterative diagonalization) are different from those stemming from TCI. A tolerance significantly below $\tau=10^{-3}$ may be desirable for two reasons: First, to avoid errors (19) larger than our mpNRG accuracy estimate of $10^{-3}$: After all, a local error (19) below the tolerance is only expected within the set of pivots that have been sampled by TCI – and even for these, the tolerance is not strictly guaranteed by the TCI routine used here (see [38], Sec. 4.3.1) , which breaks full nesting conditions. Lowering the tolerance increases the confidence that the required accuracy has been reached even outside the sampled set. The second motivation is to assess the potential of our approach for situations where more precise input data is available. We shall therefore investigate tolerances ranging from $10^{-2}$ to $10^{-5}$.
The vertex functions $\Gamma_{\mathrm{core}}(\omega,\nu,\nu^{\prime})$ and $\Gamma(\omega,\nu,\nu^{\prime})$ to be represented in QTT format here generally have prominent structures around the origin, along the frequency axes, and along the diagonals [47]. This is exemplified in Fig. 3, which shows slices of $\Gamma_{\mathrm{core}}$ at fixed bosonic frequency $\omega=0$. The inverse temperature is $\beta=2000$. We compare reference data with the QTT representation of the vertex for a TCI tolerance of $\tau=10^{-3}$. The reference was obtained by evaluating the vertex only on the two-dimensional slice shown in Fig. 3. A logarithmic color scale has been chosen to expose imperfections of the TCI approximation. Fig. 3 illustrates how QTCI represents important features of the vertex in a strongly compressed format: For a $256\!\times\!256\!\times\!$ 256 × 256 ($R=8$) frequency grid, we have ranks of $\chi=107$ for $\Gamma_{\mathrm{core}}^{\uparrow\uparrow}$ and $\chi=106$ for $\Gamma_{\mathrm{core}}^{\uparrow\downarrow}$. This translates to memory footprints reduced by factors of 92 (268 MB to 2.9 MB) and 89 (268 MB to 3.0 MB), respectively.
@@
[…]
@@
[…]
@@
We now turn to computations of the Keldysh vertex in QTT format. In contrast to the Matsubara vertex, this object gives direct access to real-frequency dynamic response functions, but is a significantly more complicated function on a continuous domain of real frequencies. Faithfully capturing its structure on a finite grid while keeping the computational cost in check is very challenging. This has been achieved in Refs. 20, 21, but requires tedious manual tuning of nonlinear grids. By contrast, our QTCI-based approach allows us to automatically capture features on different length scales on an extremely fine equidistant grid. Our grid for $\omega$ contains 0, while $\nu$ and $\nu^{\prime}$ live on a grid that is offset from 0 by half a grid spacing. An alternative choice would be to include 0 in all three grids. In QTCI, we can refine the grid until all features are represented up to a given tolerance, so that shifting the $\nu$ and $\nu^{\prime}$ grids by half a grid spacing does not make a difference. The resolution attained with QTCI is exemplified in Fig. 6, showing a QTT representation of $\Gamma_{\mathrm{core}}^{2121,\uparrow\uparrow}$ on a slice at $\omega=0$ using $R=13$ quantics bits. The TCI tolerance was set to $10^{-2}$. All panels show the same slice, but zoom in by factors of 2 moving from left to right. The rightmost panel still exhibits a sharp resolution after a 32-fold magnification. As a further illustration, Fig. 7 compares the TCI-compressed vertex (center) to the reference (left), showing the normalized error on the right. The slices are taken again at $\omega=0$ and for $u=0.5$ (top row) and $u=1.5$ (bottom row). Overall, we see that TCI resolves the core vertex to $1\%$ precision with ranks of $184$ and $187$, respectively.
@@
Figure 8 shows (a) the rank and (b) the memory size of the compressed Keldysh core vertex component $\Gamma_{\mathrm{core}}^{2121,\uparrow\uparrow}$ at temperatures $\beta=200$ and $2000$ versus the number of quantics bits $R$ in each dimension. This Keldysh component of $\Gamma_{\mathrm{core}}^{\boldsymbol{k}}$ was found to have the highest bond dimension (see App. C, Fig. 12). We set a fixed box size of $\omega_{\max}=0.65$ (cf. Fig. 6(a)) and increase the resolution with $R$. We verified that the chosen box size is large enough to capture all relevant structures within the target tolerance $\tau=10^{-2}$. At this tolerance, the rank shown in Fig. 8 saturates at $\chi=202$. This rank is again of a magnitude where a self-consistent parquet calculation in the Matsubara formalism was shown to be feasible on a single core in Ref. 28. The 16 components of the Keldysh vertex can be inferred from just 5 components using complex conjugation and crossing symmetry [50]. Nevertheless, in follow-up computations such as solving the parquet equations, these multiple Keldysh components in contrast to a single Matsubara vertex may necessitate parallelization already for $\chi\approx 200$. Multithreaded or distributed schemes will certainly be required for the most difficult case considered here ($\tau=10^{-3}$ and $\beta=2000$), which results in ranks of $\chi\approx 450$. On a different note, the QTT vertex has a vastly reduced memory footprint, as shown in Fig. 8(b): For $\beta=2000$, $\tau=10^{-2}$ and $R=10$, it requires $11.3\leavevmode\nobreak\ \mathrm{MB}$ of memory, compared to $17.1\leavevmode\nobreak\ \mathrm{GB}$ for a dense grid representation; this corresponds to a compression ratio of $1:1513$. Although real-frequency diagrammatic calculations for the SIAM are limited by runtime rather than memory [20], this paves the way for investigation of more complicated models with multiple orbitals or momentum dependence, which have prohibitive memory requirements if attempted with dense grids [51, 52, 53].
@@
Finally, we discuss how compressing each Keldysh component of the vertex separately, as was done in this work, compares to running TCI on the entire Keldysh core or full vertex, where the Keldysh components are encoded in an additional leg of a single tensor train. In both cases, spin components are compressed separately. For a fair comparison of these two approaches, recall the following: (i) TCI measures the interpolation error relative to the supremum norm of the target function (see Eq. (19)). When compressing the entire vertex, any given Keldysh component $\Gamma^{\boldsymbol{k}}$ (or $\Gamma_{\mathrm{core}}^{\boldsymbol{k}}$) should therefore be normalized by $||\Gamma^{\boldsymbol{k}}||_{\infty}$ (or $||\Gamma_{\mathrm{core}}^{\boldsymbol{k}}||_{\infty}$), i.e., with the supremum norm of the same Keldysh component $\boldsymbol{k}$. Only then does one achieve the same accuracy as in separate compressions of Keldysh components. (ii) MPO–MPO contractions, the most expensive operations in QTT-based diagrammatic calculations, scale as ($\order{R\chi^{4}}$ ) in runtime. As a preliminary investigation, we compressed the entire core vertex at $u=0.5,\beta=2000$ and $\omega_{\max}=0.65$ with a tolerance of $\tau=10^{-2}$ and $R=8$ quantics bits. The tensor leg for the Keldysh component was placed to the very left and only included the five Keldysh components not related by crossing symmetry or complex conjugation (cf. App. C).

### v3

@@ II.2 Matsubara formalism
$K(\boldsymbol{\Omega}_{p})=\!\begin{cases}&\hskip-8.5359pt\prod_{i=1}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}^{-1}\hfill\penalty 10000\ \text{if}\penalty 10000\ \prod_{i=1}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}\neq 0,\\[8.5359pt] &\hskip-8.5359pt\!-\frac{1}{2}\!\biggl[\beta+\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}^{-1}\biggr]\!\prod\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{\ell-1}\Omega_{\overline{1}\cdots\overline{i}}^{-1}\penalty 10000\ \text{if}\penalty 10000\ \exists j:\Omega_{\overline{1}\cdots\overline{j}}=0,\end{cases}$ (7)
@@
$G_{p}(\mathrm{i}\boldsymbol{\omega}_{p})=\hskip-4.30554pt\sum_{{\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}}}\hskip-4.30554ptS[\boldsymbol{\mathcal{O}}_{p}]({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}})\prod_{i=1}^{3}(\mathrm{i}{\omega}_{\bar{1}\cdots\bar{i}}-{\epsilon}_{\bar{1}\cdots\bar{i}})^{-1}{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}{.}}$ (8)
@@ II.3 Keldysh formalism
$K_{b}^{[\lambda]}(\boldsymbol{\omega}_{p},\boldsymbol{\epsilon}_{p})$ $=\prod_{j=1}^{\ell-1}\lim_{\gamma_{0}\rightarrow 0^{+}}\int_{\mathbb{R}}\mathrm{d}\omega_{\overline{1}\ldots\overline{j}}^{\prime}\,\frac{\delta_{b}(\omega_{\overline{1}\ldots\overline{j}}^{\prime},\epsilon_{\overline{1}\ldots\overline{j}})}{\omega_{\overline{1}\ldots\overline{j}}-\omega_{\overline{1}\ldots\overline{j}}^{\prime}+\mathrm{i}\gamma^{\lambda}_{0,j}}\,{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}{,}}$ (10b)
@@
[…]
@@ II.4 Symmetric improved estimators: from correlators to the vertex
is a $2\!\times\!2$ matrix acting on the $i$th Keldysh index of $G_{\mathrm{con}}$. In the Matsubara formalism, $X$ is replaced by scalar unity. In practice, the main workload in computing $\Gamma_{\mathrm{core}}$ at a given frequency $\boldsymbol{\omega}$ is the evaluation of all $2^{4}=16$ connected correlators, each comprised of $4!=24$ partial correlators. The quantities $\mathcal{K}_{1}^{r}$ and $\mathcal{K}_{2^{(\prime)}}^{r}$ can be computed using an analogous approach presented in App. B.
@@ II.5 Quantics Tensor Cross Interpolation
[…]
@@ II.6 Implementation details
[…]
@@ II.6.1 Matsubara vertices
$G_{p}(\mathrm{i}\boldsymbol{\omega}_{p})=\!\hskip-4.30554pt\sum_{{\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}}}\hskip-4.30554ptk_{\omega_{\bar{1}}\epsilon_{\bar{1}}}k_{\omega_{\bar{1}\bar{2}}\epsilon_{\bar{1}\bar{2}}}k_{\omega_{\bar{1}\bar{2}\bar{3}}\epsilon_{\bar{1}\bar{2}\bar{3}}}S[\boldsymbol{\mathcal{O}}_{p}]({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}}).$ (20)
Since $({\omega}_{\bar{1}},{\omega}_{\bar{1}\bar{2}},{\omega}_{\bar{1}\bar{2}\bar{3}})$ and $({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}})$ live on finite frequency grids, the kernels $k_{{\omega}_{\bar{1}\cdots\bar{i}}{\epsilon}_{\bar{1}\cdots\bar{i}}}$ can be viewed as matrices. A typical grid size for the spectral function peaks $({\epsilon}_{\bar{1}},{\epsilon}_{\bar{1}\bar{2}},{\epsilon}_{\bar{1}\bar{2}\bar{3}})$ is $70\times 70\times 70$, while the largest grids for Matsubara frequencies $({\omega}_{\bar{1}},{\omega}_{\bar{1}\bar{2}},{\omega}_{\bar{1}\bar{2}\bar{3}})$ used in this work have $2^{12}=4096$ points in each dimension. Thus $k_{{\omega}_{\bar{1}\cdots\bar{i}}{\epsilon}_{\bar{1}\cdots\bar{i}}}$ can be precomputed and stored for all relevant grid sizes.
@@
(i) Compressing the kernels is the more general of the two methods, in that it has a smaller memory footprint (<1<1 GB per full correlator $G(\mathrm{i}\boldsymbol{\omega})$ for $R=12,\tau=10^{-3}$). The idea is to compress the kernels $k_{\omega\epsilon}$ by exploiting their low-rank structure [39, 40]: We SVD-decompose each $k_{\omega\epsilon}$ and discard singular values below a given cutoff $S_{\mathrm{cut}}$, resulting in the approximation
@@
$\left|\sum_{\sum_{i}a_{i}>N}\hskip-8.50006ptA_{a_{1}a_{2}a_{3}}\prod_{i=1}^{3}U_{\omega_{i}a_{i}}\right|$
$\qquad\leq\sqrt{\sum_{\sum_{i}a_{i}>N}\hskip-8.50006pt|A_{a_{1}a_{2}a_{3}}|^{2}}\,\prod_{i=1}^{3}\max_{\omega_{i}}\sqrt{\sum_{a_{i}}\left|U_{\omega_{i}a_{i}}\right|^{2}}$
$\qquad=\sqrt{\sum_{\sum_{i}a_{i}>N}\hskip-8.50006pt|A_{a_{1}a_{2}a_{3}}|^{2}}\;.$ (24)
@@
In this approach, we only have two kernel contractions in each evaluation $G_{p}$, but have to store the intemediates (26) for all partial correlators. Their size grows linearly in the grid frequency grid size, i.e., as $2^{R}$ where $R$ is the number of quantics bits. For example, if $R=12$ and the PSFs live on a $70\!\times\!70\!\times\!70$ logarithmic grid (which results from $2\times 6$ decades of energy bins with 8 points per decade and discarding zeros in the PSFs), each full correlator consumes $12.9$ GB of memory. On an $R=12$ grid at $\beta=2000,\,u=0.5$, the precomputation also yields a speedup of about a factor 60. But in contrast to the compression of kernels (Eq. (21)), this speedup is independent of the TCI tolerance $\tau$. Overall, method (ii) is recommended as long as its memory demands can be met, because it evaluates correlators in a numerically exact fashion.
@@ II.6.2 Keldyh vertices
[…]
@@
Figure 2: Singular values $S_{i}$ of regular Matsubara kernels (Eq. (7), solid line) and broadened, fully retarded Keldysh kernels (Eq. (28a), dashed line). We show kernels at inverse temperatures $\beta\in\{20,200,2000\}$ and interaction $u=0.5$. The frequency grids are bosonic with $2^{12}$ points, with the Keldysh grid ranging from 0.65-0.65 to $\omega_{\max}=0.65$.
@@ III.1 Single impurity Anderson model
[…]
The vertex functions $\Gamma_{\mathrm{core}}(\omega,\nu,\nu^{\prime})$ and $\Gamma(\omega,\nu,\nu^{\prime})$ to be represented in QTT format here generally have prominent structures around the origin, along the frequency axes, and along the diagonals [47]. This is exemplified in Fig. 3, which shows slices of $\Gamma_{\mathrm{core}}$ at fixed bosonic frequency $\omega=0$. The inverse temperature is $\beta=2000$. We compare reference data with the QTT representation of the vertex for a TCI tolerance of $\tau=10^{-3}$. The reference was obtained by evaluating the vertex only on the two-dimensional slice shown in Fig. 3. A logarithmic color scale has been chosen to expose imperfections of the TCI approximation. Fig. 3 illustrates how QTCI represents important features of the vertex in a strongly compressed format: For a $256\!\times\!256\!\times\!256$ ($R=8$) frequency grid, we have ranks of $\chi=107$ for $\Gamma_{\mathrm{core}}^{\uparrow\uparrow}$ and $\chi=106$ for $\Gamma_{\mathrm{core}}^{\uparrow\downarrow}$. This translates to memory footprints reduced by factors of 92 (268 MB to 2.9 MB) and 89 (268 MB to 3.0 MB), respectively.
@@
To assess the range of applicability of our approach, we also examined how the QTT rank for $\Gamma_{\mathrm{core}}$ as well as the full vertex $\Gamma$ depends on the desired tolerance, inverse temperature $\beta$ and interaction strength $u$. The results for the $\uparrow\uparrow$ flavor and an $R=10$ grid are summarized in Fig. 5. The full and core vertices show a similar increase in rank with the TCI tolerance (Fig. 5(a)), since $\Gamma_{\mathrm{core}}$ contains precisely the complex $3$-dimensional structure of the full vertex. Consistent with previous results on random pole based Matsubara correlators (cf. Fig. 8b in Ref. 48 ), the TCI ranks increase with $\beta$, though only logarithmically, see Fig. 5(b). Finally, Fig. 5(c) shows the ranks versus the interaction strength. The key finding is that both vertices remain strongly compressible with ranks $\leq 110$ when increasing $u$ from the perturbative regime ($u\ll 1$) to very strong coupling ($u=5$). Since the $y$-axis ranges only from 85 to 110, the observed variation in ranks with $u$ does not carry much significance. Over all, Fig. 5 suggests that parquet calculations with an mpNRG vertex as input will be feasible across a wide range of parameters.
@@
According to Eq. (11), the full vertex also contains lower-dimensional contributions $K_{1}^{r}$ and $K_{2^{(\prime)}}^{r}$. In App. B, we verify that they have very small TCI ranks ($\chi\lesssim 20$ for $\tau=10^{-3}$) compared to $\Gamma_{\mathrm{core}}$ and $\Gamma$. Moreover, we focused on the $\uparrow\uparrow$ component of the vertex in the $p$-channel. In App. C, we discuss how the ranks of $\Gamma_{\mathrm{core}}$ and $\Gamma$ depend on the frequency channel and spin component.
@@
We now turn to computations of the Keldysh vertex in QTT format. In contrast to the Matsubara vertex, this object gives direct access to real-frequency dynamic response functions, but is a significantly more complicated function on a continuous domain of real frequencies. Faithfully capturing its structure on a finite grid while keeping the computational cost in check is very challenging. This has been achieved in Refs. 20, 21, but requires tedious manual tuning of nonlinear grids. The evaluation of mpNRG Keldysh vertices on nonlinear grids is discussed in App. E. Our QTCI-based approach allows us to automatically capture features on different length scales on an extremely fine equidistant grid. Our grid for $\omega$ contains 0, while $\nu$ and $\nu^{\prime}$ live on a grid that is offset from 0 by half a grid spacing. An alternative choice would be to include 0 in all three grids. In QTCI, we can refine the grid until all features are represented up to a given tolerance, so that shifting the $\nu$ and $\nu^{\prime}$ grids by half a grid spacing does not make a difference. The resolution attained with QTCI is exemplified in Fig. 6, showing a QTT representation of $\Gamma_{\mathrm{core}}^{2121,\uparrow\uparrow}$ on a slice at $\omega=0$ using $R=13$ quantics bits. The TCI tolerance was set to $10^{-2}$. All panels show the same slice, but zoom in by factors of 2 moving from left to right. The rightmost panel still exhibits a sharp resolution after a 32-fold magnification. As a further illustration, Fig. 7 compares the TCI-compressed vertex (center) to the reference (left), showing the normalized error on the right. The slices are taken again at $\omega=0$ and for $u=0.5$ (top row) and $u=1.5$ (bottom row). Overall, we see that TCI resolves the core vertex to $1\%$ precision with ranks of $184$ and $187$, respectively.
@@
Figure 8 shows (a) the rank and (b) the memory size of the compressed Keldysh core vertex component $\Gamma_{\mathrm{core}}^{2121,\uparrow\uparrow}$ at temperatures $\beta=200$ and $2000$ versus the number of quantics bits $R$ in each dimension. This Keldysh component of $\Gamma_{\mathrm{core}}^{\boldsymbol{k}}$ was found to have the highest bond dimension (see App. C, Fig. 12). We set a fixed box size of $\omega_{\max}=0.65$ (cf. Fig. 6(a)) and increase the resolution with $R$. We verified that the chosen box size is large enough to capture all relevant structures within the target tolerance $\tau=10^{-2}$. At this tolerance, the rank shown in Fig. 8 saturates at $\chi=202$. This rank is again of a magnitude where a self-consistent parquet calculation in the Matsubara formalism was shown to be feasible on a single core in Ref. 28. The 16 components of the Keldysh vertex can be inferred from just 5 components using complex conjugation and crossing symmetry [50]. Nevertheless, in follow-up computations such as solving the parquet equations, these multiple Keldysh components in contrast to a single Matsubara vertex may necessitate parallelization already for $\chi\approx 200$. Multithreaded or distributed schemes will certainly be required for the most difficult case considered here ($\tau=10^{-3}$ and $\beta=2000$), which results in ranks of $\chi\approx 450$. On a different note, the QTT vertex has a vastly reduced memory footprint, as shown in Fig. 8(b): For $\beta=2000$, $\tau=10^{-2}$ and $R=10$, it requires $11.3\penalty 10000\ \mathrm{MB}$ of memory, compared to $17.1\penalty 10000\ \mathrm{GB}$ for a dense grid representation; this corresponds to a compression ratio of $1:1513$. Although real-frequency diagrammatic calculations for the SIAM are limited by runtime rather than memory [20], this paves the way for investigation of more complicated models with multiple orbitals or momentum dependence, which have prohibitive memory requirements if attempted with dense grids [51, 52, 53].
@@
[…]